OpenStudy (anonymous):

Solve D = 3(2h – k) for h.

OpenStudy (anonymous):

Its order of operations essentially.

OpenStudy (anonymous):

A) h = D - 3k/6 B) h = D + 3k/6 C) h = D + k/6 D) h = D/3 + k

OpenStudy (anonymous):

D

OpenStudy (lopezking1):

D = 3(2h - k) 6h - k = D 6h = D + k h = (D + k)/6 Answer: h = (D + k)/6

OpenStudy (anonymous):

woah @BronxWarrior and @lopezking1 you both said different answers

OpenStudy (lopezking1):

Who did the math wrong then :o